reached at 25.6 mL of NaOH. The pH at this point is 3.59 By evaporating the solvent art the extrect, she determines that there is only 0.124 g of acid per 10.0 mLar plant extract. hat is the Ka value tor the acid? What is the pH ot the plant extract? What is the pH after the toikowing addnions of 0.1DO M Nacil: 30.0 mL. 51.2 mL, 55.0 mL What is the molecula weight of the pl ant extract? e. From the following intormation, calculate the solutility product or AgcN. Agis CN -0.07 V 0.80 V Calculate the solutility if the pH 475 and the IHCNI is 0.130 M. and the pK for HCN is 931 NT: use AG relationships between the ecuiibriumaanstant and the cel poderrtia.
Answer #1

pH = pKa + log[salt]/[acid]

given pH= 4.75, pka= 9.31 [HCN]= 0.130

4.75 = 9.31 + log [salt]/0.130

log [salt]= -5.446, [salt] = 10-5.446 = 3.5806 x 10-6

hence [AgCN]= 3.5806 x 10-6

hence Solubility, S = 3.5806 x 10-6

\Delta Go= -2.303RT log K

\Delta Go= -nFEocell

Eo cell= E(cathode)- E(anode) = 0.80 - (-0.07) = 0.87v

\Delta Go= - 1 * 96500 * 0.87 = - 83955 J

\Delta Go= -2.303RT log K

-83955 = -2.303 (8.314) (298) log K

log k= 14.713, K = 5.16 x 1014

dissociation of AgCN= Ag+ + CN-

relation betwen Ksp and solubility

Ksp = S(x+y) * xx * yy

Coefficient of Ag (x)=1

Coefficient of CN(y)=1

Ksp = (3.5806 x 10-6)2 = 7.1612 x 10 -12